OpenStudy (anonymous):

Write an equation of a line that is parallel to the line y = -4x - 5? Perpendicular? If you answer correctly, I'll select as best response and fan if not already.

OpenStudy (texaschic101):

parallel lines will have the same slope but different y intercepts In y = mx + b form, the slope is in the m position and the y intercept is in the b position. y = -4x - 5 parallel line to this is : y = -4x + 5 (or any other number besides -5) perpendicular lines have negative reciprocal slopes. All that means is " flip " the slope over and change the sign. y = -4x - 5.....slope is -4.....flip it and change the sign to find negative reciprocal. So you negative reciprocal of -4 is 1/4 (see how I flipped the slope and changed the sign) So a perpendicular line to this could be : y = 1/4x - 5 any questions ?

OpenStudy (anonymous):

no thnx :D it makes way more sense when u say it than when my teacher taught me it.